# Call Me When Necessary: LLMs can Efficiently and Faithfully Reason over Structured Environments

### v1

@@ Abstract
Large Language Models (LLMs) have shown potential in reasoning over structured environments, e.g., knowledge graphs and tables. Such tasks typically require multi-hop reasoning, i.e., match natural language utterance with instances in the environment. Previous works adopt LLMs to incrementally build a reasoning path, where LLMs either invoke tools or pick up items by step-by-step interacting with the environment. We propose Reasoning-Path-Editing (Readi), a novel framework where LLMs can efficiently and faithfully reason over structured environments. In Readi, LLMs initially generate a reasoning path given a query, and edit the path only when necessary. We instantiate the path on structured environments and provide feedback to edit the path if anything goes wrong. Experimental results on three KGQA and two TableQA datasets show the effectiveness of Readi, significantly surpassing all LLM-based methods (by Hit@1 9.1% on WebQSP, 12.4% on MQA-3H and 9.5% on WTQ), comparable with state-of-the-art fine-tuned methods (67% on CWQ and 74.7% on WebQSP) and substantially boosting the vanilla LLMs (by 14.9% on CWQ). Our code will be available upon publication.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable performance in NLP fields Ouyang et al. (2022); OpenAI (2023); Liang et al. (2023). To further unleash their reasoning ability in complex scenarios, delicate strategies are proposed to equip LLMs with human-like thought process (e.g., Chain-of-thought Wei et al., 2022) or leverage them as autonomous agents capable of planning, reflection and execution Yao et al. (2023); Liu et al. (2024b). One compelling scenario where LLMs showcase their potential is reasoning over structured environments (SEs) Jiang et al. (2023a); Sun et al. (2024). With dedicate schemas, SEs (e.g., knowledge graphs, tables) abstract real-world semantics for representing, storing, and querying data with relational structures. For instance, Freebase Bollacker et al. (2008) captures 45M entities and 3B facts over 100 domains, organized in triple patterns. The crux of successful reasoning lies in bridging the gap between natural language and the mechanism of how the SEs are represented and operated Gu et al. (2023); Li et al. (2023).
@@
While LLMs exhibit promising capabilities, their performance often falls short when faced with multi-hop reasoning involving large-scale SEs. To faithfully reason, prior works adopt an iterative way that start from certain elements (e.g., entity, relation in KG, column in Table), instantiate on SEs and then gradually expand the reasoning path while consuming just relevant portions instead of the entire environment Sun et al. (2024); Jiang et al. (2023a). This step-by-step interaction could mitigate hallucination of LLMs to some extent. However, the reasoning efficiency is sacrificed and thus hinder the practical feasibility. For a simple constraint, say “the daughter of Obama”, such methods probably require two LLM-calls to first query the relations around “Obama” and then select “father_of” from returned candidates. Moreover, at each step, the LLMs makes one choice based on history, making it prone to error propagation. Alternatively, fine-tuned methods inject environments into model parameters by tuning with human-labeled supervision. During inference, they recall schema patterns to build reasoning paths Zhang et al. (2022); Saxena et al. (2020); Luo et al. (2024); Ding et al. (2024) without interaction with SEs. This end-to-end paradigm is efficient, however, it is never ensured that the model output can be grounded on SEs. Study shows that 50% paths of RoG Luo et al. (2024) failed to yield faithful results. In addition, they heavily rely on annotations, which are difficult to obtain for large-scale SEs. Further research is thus needed to achieve efficient (i.e., less LLM-calls) and faithful (i.e., grounded on SEs) reasoning over massive SEs.
@@
Therefore, we seek to propose an interaction paradigm that leverage LLMs to support complex reasoning on large-scale SEs faithfully and efficiently. Conducting empirical study on KGQA task, we find that 46%-60% of reasoning paths initially generated by LLM can be well instantiated, which inspires us to fully exploit LLM’s intrinsic planning ability in complex reasoning. While the idea of plan-and-refine is straightforward and applied in various real world tasks Qiao et al. (2023); Collaboration (2023); Ahn et al. (2022), it’s worth noting that there exists few research on application in SEs when the initial plan encounters obstacles.
@@
In this paper, we propose Reasoning-Path-Editing (Readi), a novel framework that leverages the intrinsic planning ability of LLMs (Figure 1). In Readi, LLMs initially generate a reasoning path, which is then instantiated on SEs to facilitate faithful reasoning. Path editing is triggered only if corrections are necessary. This way, we alleviate the burden of step-by-step interaction for LLMs, resulting in improved overall efficiency. To harness the information of large-scale SEs, instead of injecting the entire static SEs into the model, we collect the reasoning log as immediate feedback which includes details such as position of stuck points, associated relations, half-way done instances, etc. This dynamic guidance refines the reasoning path more targeted and further enhances faithfulness. Our experiments on Question Answering over Knowledge Graph (KGQA) and table (TableQA) demonstrate that Readi surpasses existing solutions in terms of both efficiency and accuracy.
@@
$\bullet$ We also give detailed analysis which highlights the performance of Readi’s reasoning path generation and editing modules. Experiments demonstrate that Readi, with an average of 1.55 calls for editing, significantly reduces the number of LLM calls compared to the step-by-step interaction paradigm (which costs 4 to 8 calls). Furthermore, the reasoning log reveals that Readi exhibits characteristics akin to human thought process.
@@ 2 Related Works
Step-by-step reasoning over structured environments by LLMs. Introducing massive SEs (e.g., Freebase Bollacker et al. (2008) captures 45M entities and 3B facts over 100 domains) into LLMs context windows is not practical. Existing works break the task down to incrementally construct a reasoning path. They either treat LLMs as an agent to invoke some tools based on history states and observations  Liu et al. (2024b); Qin et al. (2023), or design specific procedures where LLMs are responsible for picking up elements on SEs Gu et al. (2023); Jiang et al. (2023a); Sun et al. (2024). These works reach faithfulness by leveraging the language processing ability of LLMs for tool or element selection. However, the performance is actually concerned with three shortcomings: 1) The iterative interaction with SEs is cumbersome, requiring quite a few LLM-calls, which is especially not efficient for complex reasoning task. 2) The greedy step-by-step decision lacks a global view of the path, making it prone to error propagation. 3) The prompts are lengthy where LLMs may lose attention of history or candidates Liu et al. (2024a). To ease these problems, we propose to directly generate a reasoning path, and edit it only when necessary, i.e., the instantiation gets stuck.
End-to-End reasoning over structured environments by fine-tuning. Fine-tuned models memorize the environments through training over annotations. They either directly generate a path and then ground it on SEs Luo et al. (2024); Huang et al. (2023b); Shu et al. (2022); Hu et al. (2022), or retrieve relevant items to build a path Zhang et al. (2022); Saxena et al. (2020); Ding et al. (2024). These end-to-end methods show efficiency with no interaction with SEs. However, they have three limitations: 1) The grounding of reasoning path only depends on model outputs, without ensuring faithfulness on SEs. To remedy this issue, they rely on a wider beam at the expense of more retrieved instances. 2) They rely heavily on annotations, which are expensive for massive environments. 3) The performance drops substantially on data unseen during training Gu et al. (2021); Huang et al. (2023a), which is common in real-world scenarios. To alleviate these problems, instead of fine-tuning, we propose to instantiate LLMs reasoning path. If anything goes wrong, we call LLMs to edit the path.
@@ 3 Task Definition
Reasoning over structured environments (SEs), e.g., question answering over knowledge graphs (KGQA) and tables (TableQA), typically requires matching a natural language question with path instances in SEs to constrain the answer. An intermediate reasoning path is a bridge between the question description and SEs. Figure 2 exemplifies some reasoning paths and their instances on KG. Note that a KG is a set of triple patterns, i.e., $\{(e,r,e^{\prime})|e,e^{\prime}\in\mathcal{E},r\in\mathcal{R}\}$, where $\mathcal{E}$ and $\mathcal{R}$ refers to the set of entity and relation, respectively.
@@
Figure 3: A running example of Readi on KGQA. An LLM initially generate an reasoning path for a question. Then, we instantiate it on KG. If anything goes wrong (the path from “France”), we collect some error messages and call an LLM to edit the path. Finally, an LLM answers the question based on the KG instances.
Following Sun et al. (2018), we model reasoning over structured environments as an retrieve-then-reason task, where we leverage LLMs to build the reasoning path given $Q$. Then we reason over the path instances the obtain the answer.
@@ 4.1 Overview
For better illustration, we adopt KGQA, a challenging scenario of multi-hop reasoning over massive environments, to showcase our concrete implementation. A running example is shown in Figure 3. Given a question and topic entities, we leverage the intrinsic planning ability of LLMs to generate initial reasoning path $P$ (Section 4.2). Then we instantiate $P$ on KG (Section 4.3). If the entire $P$ is successfully instantiated, we take the intersection of all instances then generate the answer. If any path in $P$ gets stuck, we collect error messages to guide further refinement (Section 4.4). Such process runs until $P$ is fully instantiated or a maximum edit time is reached.
@@ 4.2 Reasoning Path Generation
Inspired by Li et al. (2023), we leverage in-context learning (ICL) to generate the initial reasoning path, as shown in Figure 3(a). Given a question and $n$ topic entities, we utilize Chain-of-Thought (CoT) to generate the initial path, consisting of $n$ constraints starting from each topic entity. For the example in Figure 3, we have a two-constrained path (from “France” to constrain the bordering countries and from “Nijmegen” to airports serving it and then to countries containing these airports). Please refer to Appendix F for concrete prompts.
@@ 4.3 Reasoning Path Instantiation
For relation-binding, given a path (constraint) consisting of some NL relations (e.g., for “[France] border”, we have one relation “border” expressed in NL), we bind them with relation schemas in KG. Since a relation $\boldsymbol{r}$ in NL may have similar relation schemas in KG with analogical format and semantic meanings Li et al. (2023), we leverage BM25 and Contriever Izacard et al. (2022) to retrieve similar relations $\boldsymbol{\hat{r}}$ as candidates for $\boldsymbol{r}$. For example, we bind “border” with 3 candidate relations “adjoin, near_by, locations” in KG (the orange italicized relations in Figure 3(b). This way, we obtain candidate relations for all relations in $P$.
For path-connecting, note that the starting entity of each path is given. For the example in Figure 3(b), we need to check if there exists any path instance in KG from “France” where relations in this instance sequentially match the bound candidate relations of all relations in $P$. Specifically, if any relations in candidates “adjoin, near_by, locations” connects to “France” (the orange bold “adjoin”), we obtain entities satisfying the constraint “[France] border”, so the NL relation “border” is instantiated to “adjoin” on KG. Then, we repeat the same process for the remaining relations in $P$ to finally instantiate the whole path.
If any path in $P$ is not successfully instantiated, e.g., none of the candidate relations $\boldsymbol{\hat{r}}$ connects to the current entity, this is the “necessary” time to invoke editing (Section 4.4). If the entire $P$ is successfully instantiated or the maximum editing time is reached, we merge all KG instances by intersection to answer the question (Section 4.5). Please refer to Appendix B.1 for concrete implementation.
@@ 4.4 Reasoning Path Editing
For preparation, we collect some useful messages: 1) we show the reason of error. 2) we maintain the currently halfway-done instances. 3) relations around $\boldsymbol{\hat{e}_{err}}$ is accessible, which might be the candidates to amend the halfway-done instances. Finally, an LLM is called to edit the previous path based on $Q$ and these error messages.
@@ 4.5 QA Reasoning
Upon obtaining the merged KG instances $\mathcal{S}_{q}$, i.e., the intersection of instances for each path (constraint), we build an LLM-based reasoning module to answer the question $Q$. We concatenate $Q$ and $\mathcal{S}_{q}$ as input and prompt an LLM to generate the answer by ICL. Note that the form of $\mathcal{S}_{q}$ is a set of triple patterns $(entity,relation,entity^{\prime})$. For entities, we convert the mid in KG to the corresponding friendly name. For relations and compound nodes, we keep their original forms in KG. Specifically, we ask LLMs to pick the knowledge triples used to answer $Q$ in a CoT manner. For the example KG instances from “France” in Figure 3, we have 2 knowledge triples “(France, adjoin, compound node), (compound node, country, German)”, based on which, the LLM can reason that the country bordering France” is “German”.
@@ 5 Experiments
To comprehensively evaluate the reasoning ability of Readi on large-scale structured environments, we experiment on two complex tasks, KGQA and TableQA (implementation in Appendix A).
@@ 5.1 Datasets
We evaluate on three KGQA and two TableQA datasets. Detailed statistics are in Appendix C.
@@
MetaQA Zhang et al. (2018) includes KGQA questions from movie domain, with 3 sub-sets for 3 levels of difficulty, denoted as MQA-1H, MQA-2H and MQA-3H.
@@
We model all datasets as an information retrieval task, instead of a semantic parsing one.
@@
Table 1: QA performance (Hit@1) of Readi on KGQA datasets. Results of GPT3.5, GPT4 OpenAI (2023) and AgentBench Liu et al. (2024b) are run by ourselves, others are from the origin paper. Bold and underline fonts denotes the best and second-best for two types of methods, respectively. $\star$ denotes the overall state-of-the-art result.

### v2

@@ Abstract
Large Language Models (LLMs) have shown potential in reasoning over structured environments, e.g., knowledge graphs and tables. Such tasks typically require multi-hop reasoning, i.e., match natural language utterance with instances in the environment. Previous works adopt LLMs to incrementally build a reasoning path, where LLMs either invoke tools or pick up items by step-by-step interacting with the environment. We propose Reasoning-Path-Editing (Readi), a novel framework where LLMs can efficiently and faithfully reason over structured environments. In Readi, LLMs initially generate a reasoning path given a query, and edit the path only when necessary. We instantiate the path on structured environments and provide feedback to edit the path if anything goes wrong. Experimental results on three KGQA and two TableQA datasets show the effectiveness of Readi, significantly surpassing previous LLM-based methods (by 9.1% Hit@1 on WebQSP, 12.4% on MQA-3H and 9.5% on WTQ), comparable with state-of-the-art fine-tuned methods (67% on CWQ and 74.7% on WebQSP) and substantially boosting the vanilla LLMs (by 14.9% on CWQ). Our code will be available on https://aka.ms/readi.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable performance in NLP fields Ouyang et al. (2022); OpenAI (2023); Liang et al. (2022). To further unleash their reasoning ability in complex scenarios, delicate strategies are proposed to equip LLMs with human-like thought process (e.g., Chain-of-thought Wei et al., 2022) or leverage them as autonomous agents capable of planning, reflection and execution Yao et al. (2023); Liu et al. (2024). One compelling scenario where LLMs showcase their potential is reasoning over structured environments (SEs) Jiang et al. (2023a); Sun et al. (2024). With dedicate schemas, SEs (e.g., knowledge graphs, tables) abstract real-world semantics for representing, storing, and querying data with relational structures. For instance, Freebase Bollacker et al. (2008) captures 45M entities and 3B facts over 100 domains, organized in triple patterns. The crux of successful reasoning lies in bridging the gap between natural language and the mechanism of how the SEs are represented and operated Gu et al. (2023); Li et al. (2023).
@@
While LLMs exhibit promising capabilities, their performance often falls short when faced with multi-hop reasoning involving large-scale SEs. To faithfully reason, prior works adopt an iterative way that start from certain elements (e.g., entity, relation in KG, column in Table), instantiate on SEs and then gradually expand the reasoning path while consuming just relevant portions instead of the entire environment Sun et al. (2024); Jiang et al. (2023a); Huang et al. (2024). This step-by-step interaction could mitigate hallucination of LLMs to some extent. However, the reasoning efficiency is sacrificed and thus hinder the practical feasibility. For a simple constraint, say ‘‘the daughter of Obama’’, such methods probably require two LLM-calls to first query the relations around ‘‘Obama’’ and then select ‘‘father_of’’ from returned candidates. Moreover, at each step, LLMs make one choice based on history, making it prone to error propagation. Alternatively, fine-tuned methods inject environments into model parameters by tuning with human-labeled supervision. During inference, they recall schema patterns to build reasoning paths Zhang et al. (2022); Saxena et al. (2020); LUO et al. (2024); Ding et al. (2024); Xiong et al. (2024) without interaction with SEs. This end-to-end paradigm is efficient. However, it is never ensured that the model output can be grounded on SEs. Study shows that 50% paths of RoG LUO et al. (2024) failed to yield faithful results. In addition, they heavily rely on annotations, which are difficult to obtain for large-scale SEs. Further research is thus needed to achieve efficient (i.e., less LLM-calls) and faithful (i.e., grounded on SEs) reasoning over massive SEs.
@@
Therefore, we seek to propose an interaction paradigm that leverage LLMs to support complex reasoning on large-scale SEs faithfully and efficiently. Conducting empirical study on KGQA task, we find that 46%-60% of reasoning paths initially generated by LLMs can be well instantiated, which inspires us to fully exploit LLM’s intrinsic planning ability in complex reasoning. While the idea of plan-and-refine is straightforward and applied in various real world tasks Qiao et al. (2023); Collaboration (2023); Ichter et al. (2022), it’s worth noting that there exists few research on application in SEs when the initial plan encounters obstacles.
@@
In this paper, we propose Reasoning-Path-Editing (Readi), a novel framework that leverages the intrinsic planning ability of LLMs (Figure 1). In Readi, LLMs initially generate a reasoning path, which is then instantiated on SEs to facilitate faithful reasoning. Path editing is triggered only if corrections are necessary. This way, we alleviate the burden of step-by-step interaction for LLMs, resulting in improved overall efficiency. To harness the information of large-scale SEs, instead of injecting the entire static SEs into the model, we collect reasoning log as immediate feedback which includes details such as the position of stuck points, associated relations, half-way done instances, etc. This dynamic guidance refines the reasoning path more targeted and further enhances faithfulness. Our experiments on Question Answering over Knowledge Graphs (KGQA) and Tables (TableQA) demonstrate that Readi surpasses existing solutions in terms of both LLM-calls and accuracy.
@@
$\bullet$ We give detailed analysis which highlights the performance of Readi’s reasoning path generation and editing modules. Experiments demonstrate that Readi, with an average of 1.55 calls for editing, significantly reduces the number of LLM-calls compared to the step-by-step interaction paradigm (which costs 4 to 8 calls). Furthermore, the reasoning log reveals that Readi exhibits characteristics akin to human thought process.
@@ 2 Related Works
Step-by-step reasoning over structured environments by LLMs. Introducing massive SEs (e.g., Freebase Bollacker et al. (2008) captures 45M entities and 3B facts over 100 domains) into LLMs context windows is impractical. Existing works break the task down to incrementally construct a reasoning path. They either treat LLMs as an agent to invoke tools based on history states and observations  Liu et al. (2024); Qin et al. (2023), or design iterative procedures where LLMs are responsible for picking up items on SEs Gu et al. (2023); Jiang et al. (2023a); Sun et al. (2024). These works reach faithfulness by leveraging the reasoning ability of LLMs for tool or item selection. However, their performance is concerned with three shortcomings: 1) The iterative interaction with SEs is cumbersome, requiring quite a few LLM-calls, which is especially not efficient for complex reasoning tasks. 2) The greedy step-by-step decision lacks a global view of the path, making it prone to error propagation. 3) The accumulated prompts are lengthy where LLMs may lose attention of history or candidates. To ease these problems, we propose to directly generate a reasoning path, and edit it with feedback when the instantiation gets stuck.
End-to-End reasoning over structured environments by fine-tuning. Fine-tuned models memorize the environments through training over annotations. They either directly generate a path and then ground it on SEs LUO et al. (2024); Huang et al. (2023b); Shu et al. (2022); Hu et al. (2022), or retrieve relevant items to build a path Zhang et al. (2022); Saxena et al. (2020); Ding et al. (2024). Such end-to-end reasoning shows efficiency with no interaction with SEs. However, they have three limitations: 1) The grounding of reasoning path only depends on model outputs, without ensuring faithfulness on SEs. To remedy this issue, they rely on a wider beam at the expense of more retrieved instances. 2) They rely heavily on annotations, which are expensive for massive environments. 3) The performance drops substantially on data unseen during training Gu et al. (2021); Huang et al. (2023a), which is common in real-world scenarios. To alleviate these problems, instead of fine-tuning, we propose to instantiate LLMs reasoning path. Then, if anything goes wrong, we call LLMs to edit the path.
@@ 3 Task Definition
Reasoning over structured environments (SEs), e.g., question answering over Knowledge Graphs (KGQA) and Tables (TableQA), typically requires matching a question with instances in SEs to constrain the answer. An intermediate reasoning path is a structural representation of the question, as a bridge between the question and SEs. Figure 2 exemplifies some reasoning paths and their instances on KG. Note that a KG is a set of triple patterns, i.e., $\{(e,r,e^{\prime})|e,e^{\prime}\in\mathcal{E},r\in\mathcal{R}\}$, where $\mathcal{E}$ and $\mathcal{R}$ refers to the set of entity and relation, respectively.
@@
Figure 3: A running example of Readi on KGQA. An LLM initially generates an reasoning path for a question. Then, we instantiate it on KG. If anything goes wrong (the path from ‘‘France’’), we collect some error messages and call an LLM to edit the path. Finally, an LLM answers the question based on the KG instances.
Following Sun et al. (2018), we model reasoning over structured environments as an retrieve-then-reason task, where we leverage LLMs to build the reasoning path given $Q$. Then we reason over the path instances to obtain the answer.
@@ 4.1 Overview
For better illustration, we adopt KGQA, a challenging scenario of multi-hop reasoning over massive environments, to showcase concrete implementation (Refer to Appendix A for TableQA). A running example is in Figure 3. Given a question and topic entities, we leverage the planning ability of LLMs to generate the initial reasoning path $P$ (Section 4.2). Then we instantiate $P$ on KG (Section 4.3). If the entire $P$ is successfully instantiated, we take the intersection of all instances, then generate the answer. If any path in $P$ gets stuck, we collect error messages to guide further refinement (Section 4.4). Such process runs until $P$ is fully instantiated or a maximum edit time is reached. Please refer to Appendix G for concrete prompts of each module.
@@ 4.2 Reasoning Path Generation
Inspired by Li et al. (2023), we leverage in-context learning (ICL) to generate the initial reasoning path, as shown in Figure 3(a). Given a question and $n$ topic entities, we utilize Chain-of-Thought (CoT) to generate the initial path, consisting of $n$ constraints starting from each topic entity. For the example in Figure 3, we have a two-constrained path (from ‘‘France’’ to constrain the bordering countries and from ‘‘Nijmegen’’ to airports serving it and then to countries containing these airports).
@@ 4.3 Reasoning Path Instantiation
For relation-binding, given a path consisting of some NL relations (e.g., for ‘‘[France] border’’, we have one relation ‘‘border’’ expressed in NL), we bind them to relation schemas in KG. Since a relation $\boldsymbol{r}$ in NL may have similar relation schemas in KG with analogical format and semantic meanings Li et al. (2023), we leverage BM25 and Contriever Izacard et al. (2022) to retrieve similar relations $\boldsymbol{\hat{r}}$ as candidates for $\boldsymbol{r}$. For example, we bind ‘‘border’’ with 3 candidate relations “adjoin, near_by, locations” in KG (the orange italicized relations in Figure 3(b). This way, we obtain candidate relations for all relations in $P$.
For path-connecting, note that the starting entity of each path is given. For the example in Figure 3(b), we need to check if there exists any path instance in KG from ‘‘France’’ where relations in this instance sequentially match the bound candidate relations of all relations in $P$. Specifically, if any relations in candidates “adjoin, near_by, locations” connects to ‘‘France’’ (the orange bold “adjoin”), we obtain entities satisfying the constraint ‘‘[France] border’’, so the NL relation ‘‘border’’ is instantiated to “adjoin” on KG. Then, we repeat the same process for the remaining relations in $P$ to finally instantiate the whole path.
If any path in $P$ is not successfully instantiated, e.g., none of the candidate relations $\boldsymbol{\hat{r}}$ connects to the current entity, this is the necessary time to invoke editing (Section 4.4). If the entire $P$ is successfully instantiated or the maximum edit time is reached, we merge all KG instances by intersection to answer the question (Section 4.5). Please refer to Appendix B.1 for concrete implementation.
@@ 4.4 Reasoning Path Editing
For preparation, we collect some useful messages: 1) the reason of error. 2) the currently halfway-done instances. 3) relations around $\boldsymbol{\hat{e}_{err}}$ is accessible, which might be the candidates to amend the halfway-done instances. Finally, an LLM is called to edit the previous path based on $Q$ and these error messages by ICL.
@@ 4.5 QA Reasoning
Upon obtaining the merged KG instances $\mathcal{S}_{q}$, i.e., the intersection of instances for each path (constraint), we build an LLM-based reasoning module to answer the question $Q$. We concatenate $Q$ and $\mathcal{S}_{q}$ as input and prompt an LLM to generate the answer. Note that the form of $\mathcal{S}_{q}$ is a set of triple patterns $(entity,relation,entity^{\prime})$. For entities, we convert the mid in KG to the corresponding friendly name. For relations and compound nodes, we keep their original forms in KG. Specifically, we ask LLMs to pick the knowledge triples used to answer $Q$ in a CoT manner. For the example KG instances from ‘‘France’’ in Figure 3, we have 2 knowledge triples ‘‘(France, adjoin, compound node), (compound node, country, German)’’, based on which, the LLM can reason that ‘‘the country bordering France’’ is ‘‘German’’.
@@ 5 Experiments
To comprehensively evaluate the reasoning ability of Readi on large-scale structured environments, we experiment on two complex tasks, KGQA and TableQA. (Please refer to Appendix A for detailed implementation of TableQA).
@@ 5.1 Datasets
We evaluate on three KGQA and two TableQA datasets. Detailed statistics are in Appendix C.1.
@@
MetaQA Zhang et al. (2018) is a KGQA dataset from movie domain, with 3 levels of difficulty, denoted as MQA-1H, MQA-2H and MQA-3H.
@@
Table 1: QA performance (Hit@1) of Readi on KGQA datasets. Results of GPT3.5, GPT4 OpenAI (2023) and AgentBench Liu et al. (2024) are run by ourselves, others are from the origin paper. Bold and underline fonts denotes the best and second-best for two types of methods, respectively. $\star$ denotes the overall state-of-the-art result.